OpenStudy (anonymous):

Point A(4, 2) is translated according to the rule (x, y) -> (x + 1, y – 5) and then reflected across the y-axis. a) In which quadrant of the coordinate plane is point A located? b) what are the coordinates of translated point A'? in which quadrant of the coordinate plane is point A located? c) what are the coordinates of reflected point A''? in in which quadrant of the coordinate plane is point A located?

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

i'll try

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

In which quadrant of the coordinate plane is point A located? Quadrant 1 ----------------------------------------------------------- b) What are the coordinates of translated point A’? In which quadrant of the coordinate plane is point A’ located? A' is the point (5, -3) and it's located in quadrant 4 ------------------------------------------------------------ c) What are the coordinates of reflected point A’’? In which quadrant of the coordinate plane is point A’’ located? A'' is the point (-5, -3) and this point is located in quadrant 3 This should be correct, I'm not exactly sure. Hope this helps you. Have a nice day :)

OpenStudy (anonymous):

which equation is true for all x- and y- values in the table below? x y 1 7 2 14 3 21 4 28 y= x+ 7 y= x- 7 y= x/7 y= 7x

OpenStudy (anonymous):

y=7x

OpenStudy (anonymous):

@SimplyE please do not do direct answers. I would say y = x/7 (((c. Notice that all numbers are currently odd which will (simple tip) result in it having a / in the answer.

OpenStudy (anonymous):

how is 1/7 = 7 ?

OpenStudy (anonymous):

@StudyShark2000 please explain

OpenStudy (anonymous):

@SimplyE in the change the answer choice is a odd number. And the question contains via all numbers that it must have a /. Even if 1/7 actually equals 0.14285714285. (0.14)

OpenStudy (anonymous):

@StudyShark2000 okay still makes no sense but okay

OpenStudy (is3535):

NICE PIC

OpenStudy (is3535):

a) point A (4,2) is in the first quadrant b) translated point A' ( 5, -3) is in the fourth quadrant c) reflected point A'' ( -5,-3) is in the third quadrant ................y-axis.......... ..................|................. ..................|......A........ ---------------|----------------- x axis .......A''.......|.........A'.....

OpenStudy (anonymous):

@is3535 back off my lady lol. @SimplyE yeah it makes honestly no sense at all but it's a small tip and fairly helpful tip that our mathematics instructor/professor taught us a few years back.

OpenStudy (anonymous):

ok so the answer is c

OpenStudy (anonymous):

i mean d

OpenStudy (anonymous):

Answer would be D. I would only go with D for your particular school as I don't know if they ever taught that tip I gave you. If it does not sound familiar then just pick D as most schools will do that where if you found out a tip outside of the classroom that they did not teach they tend to say it's wrong and never test you with that particular tip.

OpenStudy (anonymous):

@darkmeadow